multiple of four consecutive natural number is 1680. find this number?
Answer by JulietG(1812)

A+B+C+D = 1680
B = A+1
C = A+2
D = A+3
Substitute the values into the original equation.
A + (A+1) + (A+2) + (A+3) = 1680
Add
4A + 6 = 1680
Subtract 6 from each side
4A = 1674
Divide each side by 4
A = 418.5
B = 419.5
C = 420.5
D = 421.5
Check by adding these together. Hmmm... these add together, but a .5 is not a natural number. Can you please check the initial question to make certain it is phrased correctly and has the right figure? The math is correct.